OpenStudy (xxkg13xx):

Are the ratios 8/12 and 10/25 proportional ? Explain. ( Note : this is not a multiple answer question)

OpenStudy (solomonzelman):

if these ratios are proportional, then must be that: \(\color{#000000 }{ \displaystyle \frac{8}{12}=\frac{10}{25} }\)

OpenStudy (solomonzelman):

If I were to determine whether a/b and c/d are proportional, then set: \(\color{#000000 }{ \displaystyle \frac{a}{b}=\frac{c}{d} }\) By cross multiplication \(\color{#000000 }{ \displaystyle c\times b=a\times d }\) and if that is true, then a/b and c/d are proportional, but if that is not true, then a/b and c/d are not proportional ratios.

jimthompson5910 (jim_thompson5910):

Here's an example showing @SolomonZelman 's formulas Question: are the two fractions 2/3 and 4/6 proportional? In other words, are they equal? 2/3 = 4/6 2*6 = 3*4 ... cross multiply 12 = 12 ... true the last equation is true. Since each equation is equivalent to one another, this means that the first equation is true. So 2/3 and 4/6 are definitely the same fraction.